particle is projected vertically upwards from ground with speed 10m/s and second particle is at the same instant from a height of 6√3m directly above the partical with speed 10 m/s at an angle of projection 30
1
Expert's answer
2021-09-12T19:01:02-0400
"x_1=10\sqrt{\frac{2(6\sqrt{3})}{9.8}}=14.56\ m\\
x_2=\frac{(10)^2\sin{60}}{9.8}=8.84\ m"